# End-to-End Visual Autonomous Parking via Control-Aided Attention

### v1

@@ IV-B2 Modular baseline study
For comparison, we evaluate a modular autonomous driving pipeline using Hybrid A* [22, 23] for planning. Obstacles are derived from BEV segmentation, ego localization uses our motion prediction model, and a PID-controlled bicycle model handles vehicle control. This baseline achieves a TSR of 59.1% with zero NTSR, but exhibits a Timeout Rate of 21.1% and Collision Rate of 12.2%, indicating that modular pipelines, while stable, are limited in handling diverse or long-range parking scenarios.
@@ IV-C Failure case study
Figure 4: Failure case study across different baselines and our method. Each row corresponds to a different experiment: (1) reproduced E2EParking trained with limited data, (2) Modular Hybrid A* baseline, (3) original E2EParking checkpoint trained with large-scale data and longer schedule, (4) CAA-Policy (Ours). Each column shows: (a) distribution of key metrics (TSR, TFR, NTSR, CR, TR) as pie charts, (b-d) the top three representative failure modes observed in that experiment. This layout allows direct comparison of overall performance and failure types across baselines and our method. Note that Trajectories are color-coded according to time sequence. For the Modular Hybrid A* baseline, red trajectories indicate the planned GT trajectory, which the controller struggles to follow on consecutive turns.
@@
Modular Hybrid A*: As shown in the second row of Fig. 4, classical planning pipelines using Hybrid A* generally produce accurate ego and target estimates. The main factors behind residual failures are: (1) Compounding perception errors: minor inaccuracies in object detection or free-space estimation can misguide trajectory planning; (2) Online planning limitations: computing feasible paths entirely online in complex environments can be computationally intensive, limiting the update frequency of perception inputs; (3) Controller-induced tracking errors: executing planned trajectories via a separate controller (e.g., PID) can introduce deviations that accumulate over longer horizons. This is particularly noticeable in scenarios with consecutive sharp turns, where even small path errors can cause the vehicle to deviate significantly from the planned trajectory. Although Hybrid A* demonstrates robust planning capabilities, suboptimal or infeasible paths in these situations are challenging for controllers to track accurately. This motivates the design of CAA-Policy, which directly generates control commands rather than relying on complex path planning. By doing so, it reduces accumulated tracking errors.
@@ IV-D Ablation Study
We analyze the contribution of each key module in CAA-Policy - Control-Aided Attention (CAA), Target Tokenization Module (TTM), and Waypoint Prediction Head (Waypoints) - and their combinations to quantify complementary benefits and overall performance gains shown in Table III.
IV-D1 CAA module
The CAA module refines perception features to emphasize control-relevant regions. Applied alone, it performs worse than the baseline (18.7% TSR vs. 26.8%), as the backbone alone may provide insufficient guidance, causing reinforcement of incorrect attention patterns. This results in higher non-target success (20.3%) and a large collision rate (32.8%), highlighting that CAA alone cannot reliably guide the model without complementary trajectory information.
@@ IV-D2 Target Tokenization Module (TTM)
The TTM module raises TSR from 26.8% (baseline E2EParking) to 74.0%. By explicitly encoding the target parking slot, it clarifies the intended goal for the agent, stabilizing control and reducing ambiguous behaviors. It also mitigates drifting into non-target slots, with NTSR dropping from 19.9% to 0.3%. Nevertheless, some failures remain, as reflected in the collision rate (13.5%) and timeout rate (9.9%), indicating that TTM further benefits from complementary modules providing trajectory guidance.
IV-D3 Waypoint module
Applying the Waypoint module alone, the model achieves a TSR of 76.0%, a substantial improvement over the reproduced baseline (26.8%) and comparable to the original E2EParking trained on large-scale data, while reducing the Collision Rate to 9.1%. This demonstrates that explicit intermediate waypoints provide strong spatial guidance, constraining the agent’s trajectory and reducing unintended deviations. However, the Timeout Rate remains relatively high (12.2%).
IV-D4 Module Synergy
We investigate the synergy between TTM, CAA, and Waypoints (Table III). Removing any single module while keeping the other two yields TSR values around 80–81%, indicating complementary contributions. However, all partial combinations fall short of the full CAA-Policy, which achieves 87.5% TSR — roughly 6–7% higher than any ablation — showing that jointly integrating TTM, CAA, and Waypoints maximizes performance and reduces collisions, timeouts, and non-target successes.
@@ VI Conclusion
In this work, we present CAA-Policy, an end-to-end autonomous parking framework that integrates Control-Aided Attention (CAA) for task-specific perception refinement, a Target Tokenization Module (TTM) for explicit target position representation, and a Waypoint Prediction auxiliary task to enforce temporal consistency. By explicitly predicting both vehicle motion and target location, and leveraging a short-horizon vehicle motion prediction model, CAA-Policy achieves robust and accurate tracking of ego and target positions in complex environments. Extensive experiments in CARLA demonstrate that each module contributes positively to performance, and their combination leads to improved trajectory stability, control accuracy, and overall success rates. These results highlight the effectiveness of combining perception refinement, temporal reasoning, and explicit motion supervision in end-to-end autonomous driving systems.
@@
In the supplementary material, we provide additional visualizations and analyses including: (1) a comparison of different target tracking strategies, (2) an evaluation of computation cost between end-to-end and modular baselines, and (3) a robustness study under perturbations in vehicle speed and target noise.
-A Target Tracking comparison
We evaluate approaches for tracking both ego and target positions: (1) a segmentation-based method [13], (2) a motion prediction model using a single historical frame, and (3) a motion prediction model leveraging multiple historical frames.
TABLE I: Target tracking comparison using Average Displacement Error (ADE) and variance. Inference time is negligible.
@@
Dynamics (single-frame) 0.403 0.020

### v2

@@ IV-B2 Modular baseline study
For comparison, we evaluate a modular autonomous driving pipeline using Hybrid A* [dolgov2008practical, kurzer2016path] for planning. Obstacles are derived from BEV segmentation, ego localization uses our motion prediction model, and a PID-controlled bicycle model handles vehicle control. This baseline achieves a TSR of 59.1% with zero NTSR, but exhibits a TR of 21.1% and CR of 12.2%, indicating that modular pipelines, while stable, are limited in handling diverse or long-range parking scenarios.
@@ IV-C Failure case study
Figure 4: Failure case study across different baselines and our method. Each row corresponds to a different experiment: (1) reproduced E2EParking trained with limited data, (2) Modular Hybrid A* baseline, (3) original E2EParking checkpoint trained with large-scale data and longer schedule, (4) CAA-Policy (Ours). Each column shows: (a) distribution of key metrics (TSR, TFR, NTSR, CR, TR) and minor errors (e.g., out-of-bound) as pie charts. (b-d) the top three representative failure modes observed in that experiment. This layout allows direct comparison of overall performance and failure types across baselines and our method. Note that Trajectories are color-coded according to time sequence. For the Modular Hybrid A* baseline, red trajectories indicate the planned GT trajectory, which the controller struggles to follow on consecutive turns.
@@
Modular Hybrid A*: As shown in the second row of Fig. 4, despite the low NSTR suggesting that Hybrid A* [dolgov2008practical, kurzer2016path] can plan reliably toward the correct target, it frequently fails due to time run-outs and collisions. These failures primarily stem from three factors: (1) Compounding perception errors: small inaccuracies in object detection or free-space estimation can misguide trajectory planning; (2) Computation overhead: generating feasible paths to distant targets is computationally expensive, limiting the update frequency of perception inputs; (3) Controller-induced tracking errors: executing trajectories with a separate controller (e.g., PID) introduces deviations that accumulate over long horizons. This issue is especially pronounced in scenarios with consecutive sharp turns, where even minor path errors can cause large deviations. It is also worth noting that our evaluation directly integrates [dolgov2008practical, kurzer2016path] into the modular framework without extensive engineering of the number of planning and execution steps or manual tuning of search-tree expansion parameters—which could substantially improve real-world performance.
@@ IV-D Ablation Study
We analyze the contribution of each key module in DL-based Motion Prediction (MotionPred), Target Tokenization Module (TTM), and Waypoint Prediction Head (WayPts), CAA-Policy - Control-Aided Attention (CAA)- and their combinations to quantify complementary benefits and overall performance gains shown in Table III.
IV-D1 DL-based Motion Prediction (MotionPred)
The failure analysis of both the original E2EParking and our reproduced version indicates that parking to an incorrect target is a critical issue. As shown, the introduced DL-based motion prediction module significantly reduce the NTSR and improve TSR. Given its essential role, this module is included as a default component in all remaining variants of our method.
@@ IV-D2 Target Tokenization Module (TTM)
The TTM module raises TSR from 26.8% (baseline with our motion prediction) to 74.0%. By explicitly encoding the target parking slot, it clarifies the agent’s goal, stabilizing control and reducing ambiguous behaviors. It also lowers NTSR from 19.9% to 0.3%. Some failures remain—collision rate at 13.5% and timeout rate at 9.9%—showing TTM still benefits from complementary trajectory guidance modules.
IV-D3 Waypoint module (WayPts)
When applying only the Waypoint module, the model achieves a TSR of 76.0%, representing a substantial improvement over the reproduced baseline and matching the performance of the E2EParking checkpoint trained on large-scale data, while reducing the CR to 9.1%. This indicates that explicit intermediate waypoints provide strong spatial guidance, constraining the agent’s trajectory and reducing unintended deviations. However, the Timeout Rate remains relatively high (12.2%).
IV-D4 CAA module
The CAA module refines perception features by emphasizing control-relevant regions. As shown in the ablation study, combining this module with either the TTM module, the Waypoint module, or both consistently improves performance. Since the other modules already elevate the overall system to a high baseline, achieving an additional gain of over 5% in success rate is particularly valuable for meeting the demands of precision parking. We also observe that replacing CAA with CBAM [woo2018cbam] does not lead to notable improvements. This confirms the contribution of the control-sensitive supervison we introduced in CAA.
IV-D5 Module Synergy
We investigate the synergy between TTM, Waypoints and CAA. Removing any single module while keeping the other two yields TSR values around 80–81%, indicating complementary contributions. However, all partial combinations fall short of the full CAA-Policy, which achieves 87.5% TSR — roughly 6–7% higher than any ablation — showing that jointly integrating TTM, Waypoints, and CAA maximizes performance and reduces collisions, timeouts, and non-target successes.
@@ VI Conclusion
In this work, we present CAA-Policy, an end-to-end autonomous parking framework that integrates Control-Aided Attention (CAA) to couple perception with control demands, together with enhancements from a learnable motion prediction module, a Target Tokenization Module (TTM), and an auxiliary waypoint prediction task. Experiments in CARLA show that each module improves performance, while their combination yields substantial gains in trajectory stability, control accuracy, and success rate, demonstrating the effectiveness of incorporating modular designs into an end-to-end learnable system.
@@
In the supplementary material, we provide additional visualizations and analyses, including: (1) a comparison of different target tracking strategies, (2) an evaluation of computation cost between end-to-end and modular baselines, and (3) a robustness study under perturbations in vehicle speed and target noise.
-A Motion Prediction comparison
We evaluate approaches for tracking both ego and target positions: (1) a segmentation-based method [yang2024e2e], (2) a motion prediction model using a single historical frame, and (3) a motion prediction model leveraging multiple historical frames.
TABLE I: Target tracking comparison using Average Displacement Error (ADE) and variance. Inference time is negligible. Note that MP. is short for motion prediction.
@@
Learnable MP. (single-frame) 0.403 0.020